OpenStudy (bbb911):

First estimate and multiply for exact answer 6 3/5 X 5 4/11 X 6

OpenStudy (anonymous):

Im a bit confused in the way it is written.

OpenStudy (anonymous):

6 3/5*5 4/11*6?

OpenStudy (anonymous):

estimate : 7 x 5 x 6 = 30 x 6 = 210 exact answer : 6 3/5 x 5 4/11 x 6 (change mixed numbers to improper fractions) (33/5 x 59/11) x 6/1= 1947/55 x 6/1= 11682/55 = 212 22/55 = 212 2/5

OpenStudy (bbb911):

okay yea :) i got 210 for the estimate i just got lost on the way thank you so much :)

OpenStudy (anonymous):

your welcome :)